# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v4

@@ Abstract
Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence. While large language models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence. Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer little guidance for incremental progress. To address this gap, this paper introduces the artificial general intelligence testbed (AGITB), a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction. AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing. Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles. While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence.
@@ 1 Introduction
Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of human cognition’s flexible, robust intelligence. Marcus and Davis (2020) observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. This surface-level competence often obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI). Achieving AGI entails the ability to generalize, reason, and adapt across a wide range of domains in a goal-directed and context-sensitive manner.
Recent progress in AI has led to growing speculation that AGI may be approaching. However, such claims remain difficult to substantiate without a rigorous and informative evaluative framework. Assessing progress toward AGI requires more than task-specific benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing broad cognitive capabilities.
Various attempts have been made to define such tools, the most iconic being the Turing Test proposed by Turing (1950). However, existing benchmarks fail to meet the needs of AGI research and lack gradual resolution, interpretability, and automation. They typically assess surface-level competence rather than underlying mechanisms of generalization. In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial systems.
@@ 2 Background
Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics. This is partly due to the absence of a universally accepted AGI benchmark. Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified. However, such specialized benchmarks favor narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence. Ironically, some of these benchmarks are now so specialized that humans no longer reliably outperform machines on them.
@@ 2.1 A Test That Humans Pass but Machines Fail
A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity grounded in biologically inspired processing that current artificial systems cannot replicate. The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets. The second strategy suggests that a viable path toward AGI may involve more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the level of signal processing. While Turing was right to suggest that communicative behavior could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
@@ 3.1 Components
AGITB consists of two principal components: the Cortex, which represents the AGI system under evaluation, and the Input, which encodes the data samples presented to the cortex over time. The cortex is tasked with maintaining an evolving internal state and using it to generate predictions of future inputs based on the observed history of prior signals.
Each input is a binary-encoded sample that simulates information received from (virtual) sensors and actuators. These samples are composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal. Thus, an input vector represents a snapshot of multiple parallel signals at a given moment. For example, a 10-bit input might encode visual information from a $2\times 3$ pixel array and four bits of auditory or tactile data. While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences. Although spatial and temporal dimensions are orthogonal, their interaction yields structured patterns that the cortex must detect and exploit.
@@ 3.2 Operation
The testbed presents the cortex with a sequence of inputs over time. At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input sequence and using that understanding to make accurate future predictions. In cases where predictions are not exact, proximity to the correct value is considered, which indicates the model’s capacity for generalization and adaptation.
@@
To execute AGITB, users must specify one parameter: the pattern period. This value determines the length of the repeating sequence in the input stream that the cortex is expected to recognize and adapt to. Longer pattern periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities. However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity. As such, the pattern period should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
@@ 3.3 The 12 Essential Tests
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner and Kistler (2002) showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning. Consequently, AGITB permits only those sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
@@
Assertion: The cortex can successfully adapt to input sequences that repeat with the user-specified pattern period.
Assertion: The cortex can also adapt to input sequences with a pattern period longer than the specified value.
This test measures a model’s ability to generalize across temporal scales. Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of exact periodicity.
@@
Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be predictable.
Cognitive systems exhibit adaptation limits due to finite resources or saturation. This test evaluates whether such limits emerge over time, even in the presence of learnable input.
Test 9 – Unsupervised
@@ 3.4 Memorization
To prevent models from using brute-force memorization, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints—both in time and memory. In AGITB, a task typically involves predicting a sequence of $n$ binary inputs, each consisting of $k$ bits. This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps. This constraint significantly narrows the set of valid input sequences. For non-cyclic sequences, the number of valid configurations is given by $|S\textquoteright|=F_{n+2}^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
For cyclic sequences—where the pattern repeats and the first input must also satisfy the refractory constraint relative to the last—the valid space is further reduced to $|S^{\prime\prime}|=L_{n}^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
@@
Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality. Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization quickly become computationally intractable. The exponential complexity highlights the necessity for models that rely on generalizable, pattern-based learning mechanisms rather than rote memorization.
@@ 3.5 Performance
AGITB yields meaningful insights only when developers rigorously align their systems with its core assumptions. Misinterpreting elements such as the “empty” initial state can result in unreliable or misleading outcomes.
@@ 3.5.1 Human Performance
AGITB assumes that humans inherently satisfy all 12 tests, as the biological architecture of the cortex naturally supports low-level binary signal processing. Some of these mechanisms also manifest at the cognitive level, where their effects can be observed through conscious reasoning and introspection.
At present, predicting the next binary input in a temporal sequence remains the only practical method for directly evaluating human performance on AGITB at the cognitive level. However, the inability to compare internal cortical states makes Tests 2 (Bias), 3 (Determinism), 4 (Sensitivity), and 5 (Time) untestable in humans. In addition, Test 6 (Refractory Period) does not translate meaningfully to the cognitive level, as consecutive activations (e.g., binary ones) can be valid and interpretable in conscious reasoning.
@@
Tests 9 (Unsupervised), 10 (Knowledge), and 11 (Unobservability) involve comparisons across multiple individuals. However, a single person may internally simulate or approximate the alternate perspectives required, effectively fulfilling these roles in parallel with the primary task.
In general, once the task is understood, adult humans can reliably pass the applicable AGITB tests, so long as the combination of pattern period and input dimensionality remains within natural cognitive limits.
@@ 3.5.2 Classical Symbolic Programs
In classical AI systems, a model’s initial state is inseparable from its program, which encodes fixed logic and predetermined responses. Consequently, such systems cannot begin from an unbiased or knowledge-free state, as their behavior is entirely dictated by human-authored instructions. AGITB’s first test, however, introduces a paradox: the system must initialize without any such program—that is, it must begin in a truly neutral state, free from prior knowledge or assumptions. In the language of Searle’s Chinese Room Argument, the ledger must be blank—containing no predefined instructions for interpreting Chinese symbols. Yet classical systems require a program to function; absent such instructions, they halt. The paradox lies in the fact that if such systems adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem.
AGITB’s initial test thus formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules. To enforce this condition reliably, the model comparison operator must be implemented with rigor, ensuring that all relevant internal properties of the system are considered. Any structural discrepancy or hidden initialization may conceal bias and undermine the integrity of the test.
@@ 3.5.3 Artificial Neural Network Performance
Unlike classical symbolic programs, artificial neural networks do not encounter the same initialization paradox. Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their output remains neutral (for example, when an activation function returns zero).
@@ 3.5.4 Large Language Model Performance
Although built on ANN architectures, LLMs are not suited for evaluation at the signal-processing level. While their internal weights and parameters are technically accessible, AGITB’s first test requires models to begin in an unbiased, knowledge-free state—an assumption fundamentally incompatible with the nature of pre-trained systems. LLMs cannot generate meaningful predictions without pretraining, making low-level evaluation impractical and uninformative.
As with human participants, large language models (LLMs) must be evaluated at a higher cognitive level, beyond the binary signal-processing domain that AGITB was originally designed to assess. To explore whether LLMs exhibit AGI-like behavior on Tests 7 through 12—as done with humans—we manually evaluated ChatGPT-4o using a structured prompt (see Appendix). The prompt required ChatGPT to simulate two distinct models, enabling the assessment of multi-agent dynamics relevant to Tests 9, 10, and 11.
Beginning with Test 7 (Temporal Flexibility), ChatGPT-4o successfully adapted to basic period-2 patterns but failed with slightly more complex period-3 sequences, such as alternating between 000001, 000010, and 000100. It often defaulted to repeating an incorrect prior input or generated seemingly arbitrary outputs. Due to this fundamental limitation, further testing was discontinued.
@@ 3.6 Remarks
AGITB evaluates a model’s predictive capabilities following exposure to sequences of either structured or random inputs. Random input sequences with arbitrary internal correlations are used to minimize reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed by Zador et al. (2023), where cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT. AGITB’s “all-tests-must-pass” standard enforces a rigorous, biologically motivated foundation for evaluating AGI, grounded in the same principles that underlie natural intelligence.

### v5

@@ Abstract
Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence. While large language and reasoning models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence. Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer no guidance. The artificial general intelligence testbed (AGITB) is a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction. AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing. Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles. While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence. A reference implementation of AGITB is available on GitHub.
@@ 1 Introduction
Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of flexible, robust human intelligence. Marcus and Davis (2020) observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. Mitchell (2025) further summarized why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted. The surface-level competence obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI).
The progress in AI has led to growing speculation that AGI is near. However, such claims remain difficult to substantiate without a rigorous and informative metric. Assessing progress toward AGI requires more than specialised benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing (1950). However, no existing metric fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability. They typically assess surface-level competence rather than underlying mechanisms of generalization. In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial intelligence systems.
@@ 2 Background
Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics. This is partly due to the absence of a universally accepted AGI benchmark. Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified. However, such specialized benchmarks favor narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence. Ironically, some of these benchmarks are now so specialized that humans have difficulty with them.
@@ 2.1 A Test That Humans Pass but Machines Fail
A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity that current artificial systems cannot replicate. The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets. The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level. While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
@@ 3.1 Components
AGITB requires the user to define two components: the cortex, representing the AGI system under evaluation, and the input, which encodes the data samples delivered to the cortex over time. The cortex is treated as a black box that predicts future inputs based on the observed history of prior signals.
Each input consists of binary-encoded data from (virtual) sensors and actuators. It is composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal. An input represents a snapshot of multiple parallel signals at a single point in time, as illustrated in Figure 1a. While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences, as shown in Figure 1b. Although the spatial and temporal dimensions are orthogonal, their interaction produces structured patterns to which the cortex must adapt.
Figure 1: (a) Example of a 2×32$\times$ 32 × 3-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
(b) Example of an input sequence illustrating a temporal pattern with a period of 4.
@@ 3.2 Operation
The testbed presents the cortex with a stream of inputs over time. At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions. In cases where predictions are not exact, proximity to the correct value is considered, which indicates the model’s capacity for generalization and adaptation.
Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps. At the previous step ($t-1$), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data. Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
@@
To execute AGITB, users must specify one parameter: the pattern period. This value determines the length of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to. Longer pattern periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities. However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity. As such, the pattern period should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
@@ 3.3 The 12 Essential Tests
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner and Kistler (2002) showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning. Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
@@
Assertion: The cortex must be able to adapt to temporal patterns with the user-specified period.
Assertion: The cortex must also be capable of adapting to temporal patterns with periods longer than the user-specified value.
This test evaluates a model’s capacity to generalize across temporal scales. Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity. The purpose of this test is to identify models that are overly specialized to a single temporal period rather than capable of robust generalization across multiple timescales.
@@
Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be learnable.
Cognitive systems inevitably reach adaptation limits as their finite resources become saturated. This test assesses whether such limits emerge over time when the cortex is exposed to input sequences that would otherwise be learnable.
Test 9 – Entropy
@@ 3.4 Memorization
To prevent models from using brute-force memorization, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints—both in time and memory. In AGITB, a task typically involves predicting a temporal sequence of $n$ binary inputs, each consisting of $k$ bits. This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
@@
AGITB further reduces this space by requiring sequences to be cyclic, meaning that the first input must also satisfy the refractory constraint relative to the last input in the sequence. The number of distinct cyclic temporal input sequences with refractory constraint is given by $|S^{\prime\prime}|=(L_{n})^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
Consider an illustrative example of a task with 3-bit inputs and a pattern period of $n=5$. The total number of unconstrained sequences is $|S|=2^{15}=32.768$. Applying the refractory constraint reduces the valid set to $|S\textquoteright|=13^{3}=2.197$ sequences. When the cyclic condition is added, the space shrinks further to $|S^{\prime\prime}|=11^{3}=1331$.
Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality. Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization are computationally intractable. The exponential complexity highlights the necessity for models that rely on generalizable, pattern-based learning mechanisms rather than rote memorization.
@@ 3.5 Performance
AGITB can yield meaningful insights only when developers adhere faithfully to its core requirements. Misinterpreting elements such as the “empty” initial state can lead to unreliable or misleading results.
@@ 3.5.1 Human Performance
The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans. However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most of the 12 tests. Tests 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
@@
AGITB includes the refractory period requirement (Test 6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution. Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection. Temporal flexibility (Test 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations. The property of stagnation (Test 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge. The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event. Because humans learn different topics at varying speeds, the rationale for the entropy (Test 9) and the salience requirement (Test 10) is self-evident. The inability to directly inspect human mental states is captured in the unobservability requirement (Test 11). Finally, the generalization test (Test 12) requires what humans naturally excel at: generalizing when faced with new but distorted input.
@@ 3.5.2 Classical Symbolic Programs
In classical AI systems, a model’s initial state is inseparable from its program, which encodes fixed logic and predetermined responses. Consequently, such systems cannot begin from an unbiased or knowledge-free state, as their behavior is entirely dictated by human-authored instructions. AGITB’s first test, however, introduces a paradox: the system must initialize without any such program—that is, it must begin in a truly neutral state, free from prior knowledge or assumptions. If they adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem. In the language of Searle’s Chinese Room Argument, the ledger must be blank—containing no predefined instructions for interpreting Chinese symbols. Yet classical systems require a program to function; absent such instructions, they halt.
AGITB’s initial test formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules. To enforce this condition reliably, the model comparison must be rigorous, ensuring that all relevant internal properties of the system are examined. Any structural discrepancy or hidden initialization could conceal bias and compromise the integrity of the test.
@@ 3.5.3 Artificial Neural Network Performance
Unlike classical symbolic programs, artificial neural networks do not encounter the same initialization paradox. Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their input remains neutral (for example, when an activation function returns zero).
@@ 3.5.4 Large Language Model Performance
Because LLM is essentially an ANN, it fails the genesis test. This failure immediately indicates that LLMs cannot learn entirely on their own and are inherently biased toward the language and data they were trained on. Although the internal state of an LLM is readily accessible and could, in principle, be used to perform other tests, such evaluations are meaningless until the first requirement is satisfied.
Consequently, LLMs cannot acquire new knowledge—let alone learn a new language—solely from prompts, nor can they develop the kind of grounded understanding of the world that humans possess. This implication also applies to large reasoning models, which are themselves built upon LLM architectures.
@@ 3.6 Remarks
AGITB evaluates a model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs. Random input sequences with arbitrary internal correlations are used to minimize reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador et al. (2023), where cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT. AGITB’s “all-tests-must-pass” standard enforces a rigorous, biologically motivated foundation for evaluating AGI, grounded in the same principles that underlie natural intelligence.